# Disturbance Compensation for Safe Kinematic Control of Robotic Systems with Closed Architecture

### v1

@@ VI Conclusion and future work
This paper presents a unified framework that improves tracking performance and ensures safe kinematic control for closed-architecture robotic systems. The proposed method integrates seamlessly into existing cascade control structures as an add-on, enabling robust and safe control under uncertain dynamics, external disturbances, and unknown inner-loop controllers. Stability and safety are theoretically guaranteed, and real-time validation is conducted on the PUMA 500 industrial robot. Future work will extend the framework to visual servoing systems, addressing uncertainties in the vision pipeline.

### v2

@@ VII Conclusion and future work
A unified framework that enhances tracking performance and ensures safe kinematic control for robotic systems with closed architecture is proposed in this paper. Compared to the state-of-the-art method in [ahanda2022adaptive], which is used solely for tracking, our approach introduces safety features. The proposed approach can be easily integrated into the existing cascade control framework as an add-on, providing robust and safe control in the presence of various uncertainties, including uncertain dynamics, external disturbances, and an uncertain inner-loop controller. Theoretical results of stability analysis and provable safety guarantees are presented. The framework has been validated in real-time on the PUMA 500 industrial robot.
In future work, we plan to extend our approach to a visual servoing control system to address additional challenges, particularly when uncertainty in the vision system needs to be considered.
Appendix A Proof of the Bounded Series in (51)
The sum of the infinite series in (51) can be expressed as
$\begin{array}[]{r@{}l}\sum\limits_{k=1}^{\infty}p(k)=&r_{i}+1+\sum\limits_{\iota=1}^{r_{i}+1}\dfrac{1}{(\iota-1)!}(1-\omega_{o_{i}})^{\iota-1}\\ &\cdot\sum\limits_{k=1}^{\infty}\left(\prod\limits_{j=-\iota+1}^{-1}(k+j)\right)\omega_{o_{i}}^{k-\iota},\end{array}$ (60)
where the finite sum on the right-hand side is clearly bounded. The key step is to show that the second infinite sum converges. Note that since $j<0$, $k>0$, $\iota\geq 1$ and $\omega_{o_{i}}>0$, we can write
$\sum\limits_{k=1}^{\infty}\left(\prod\limits_{j=-\iota+1}^{-1}(k+j)\right)\omega_{o_{i}}^{k-\iota}\leq\sum\limits_{k=1}^{\infty}k^{\iota-1}\omega_{o_{i}}^{k-\iota}.$ (61)
Applying the ratio test, we have
$\lim_{k\to\infty}\frac{(k+1)^{\iota-1}\omega_{o_{i}}^{k+1-\iota}}{k^{\iota-1}\omega_{o_{i}}^{k-\iota}}=\lim_{k\to\infty}\left(\frac{k+1}{k}\right)^{\iota-1}\omega_{o_{i}}=\omega_{o_{i}}.$ (62)
Since $\omega_{o_{i}}<1$ in the discrete-time domain, the infinite series $\sum\limits_{k=1}^{\infty}k^{\iota-1}\omega_{o_{i}}^{k-\iota}$ converges. This completes the proof.
Appendix B Implementation Details of DOB-CBF [dacs2022robust] for Comparison
The DOB proposed in [dacs2022robust] is designed for a system with relative degree one. In other words, it estimates the effect of disturbances on the dynamics of $h$, i.e.,
$\underbrace{\ddot{h}(x)}_{\dot{h}^{{}^{\prime}}(x)}=\underbrace{L_{f}^{2}h(x)+L_{g}L_{f}h(x)u}_{a_{e}(x,u)}+\underbrace{L_{g}L_{f}h(x)f}_{b_{e}(x,f)},$ (63)
where $a_{e}(x,u)$ is known and $b_{e}(x,f)$ needs to be estimated.
In this paper, $\ddot{h}$ is expressed as follows:
$\begin{array}[]{ll}\ddot{h}&=\frac{d}{dt}\left(\frac{\partial\zeta}{\partial q}\right)\dot{q}+\frac{\partial\zeta}{\partial q}\Bigl(-\bar{M}^{-1}(\bar{C}\dot{q}+\bar{G}-\bar{K}_{d}\dot{q}_{d})+f)\Bigr)\\ &=\underbrace{\frac{d}{dt}\left(\frac{\partial\zeta}{\partial q}\right)\dot{q}+\frac{\partial\zeta}{\partial q}\Bigl(-\bar{M}^{-1}(\bar{C}\dot{q}+\bar{G}-\bar{K}_{d}\dot{q}_{d})\Bigr)}_{a_{e}(x,u)}+\underbrace{\frac{\partial\zeta}{\partial q}f}_{b_{e}(x,f)}.\end{array}$ (64)
Using the DOB-CBF approach, the unknown term to be estimated is $b_{e}(x,f)=\frac{\partial\zeta}{\partial q}f$. A DOB is designed as follows:
$\begin{cases}\begin{aligned} \hat{b}_{e}&=k_{b}\dot{h}-\chi,\\ \dot{\chi}&=k_{b}(a_{e}(x,u)+\hat{b}_{e})\\ &=k_{b}(a_{e}(x,u)+k_{b}\dot{h}-\chi),\end{aligned}\end{cases}$ (65)
where $k_{b}$ is the observer gain, $\dot{h}=\left(\frac{\partial\zeta}{\partial q}\right)\dot{q}$, and $\zeta$, $q$ and $\dot{q}$ are obtained from the measurements. Suppose $b_{e}(x,f)$ is differentiable with respect to $t$, with an upper bound on its rate of change, $|\dot{b}_{e}(x,f)|\leq b_{h}$. Then, the steady-state estimation error bound of $\hat{b}_{e}$ is given by $\frac{b_{h}}{k_{b}}$.